OpenStudy (anonymous):

If the sides of a square are lengthened by 7 cm, the area becomes 121 cm^2. Find the length of a side of the original square

OpenStudy (anonymous):

solve \[(x+7)^2=121\] \[x+7=11\] \[x=11-7=4\]

OpenStudy (anonymous):

x = length of a side of the original square Lengthen the sides by 7 cm area = 121 cm² (x+7)² = 121 x+7 = √121 x+7 = 11 x = 4 original square was 4 cm by 4 cm